OpenStudy (anonymous):

A person opens his wallet and finds that all the bills are $1 bills, $5 bills, or $10 bills. There are 43 bills in this person's wallet. There are 8 more $5 bills than $1 bills. If the total amount of bills is $278, How many $10 bills does this person have? (establish a system of three equations)

OpenStudy (mertsj):

x=ones y=fives z=tens x+y+z=43 y=x+8 x+5y+10z=278

OpenStudy (mertsj):

Solve the system